The Weidmans want to save $25,000 in 3 years for a down payment on a house. If they make monthly deposits in an account paying 12%, compounded monthly, what is the size of the payments that are required to meet their goal? (Round your answer to the nearest cent.)
Answer #1

Given that,

amount required in t = 3 years is FV = $25000

interest rate r = 12% compounded monthly

so, size of monthly payment can be calculated using formula,

monthly payment PMT = FV*(r/n)/((1+r/n)^(n*t) - 1) = 25000*(0.12/12)/((1+0.12/12)^(12*3) - 1) = $580.36

So, monthly payments required to meet the goal is $580.36